12)A data set includes data from student evaluations of courses. The summary statistics are n = 84, x = 4.18, S = 0.63. Use a 0.(Reject) (sufficient) (is)

(Fail to reject) (Not sufficient) (is not)
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 4.25

Ha : \mu \neq 4.25

Test statistic (t)

= (\bar x - \mu) / s / \sqrtn

= (4.18 - 4.25) /0.63 / \sqrt84

Test statistic = -1.018

2 * P(z < -1.018) = 2 * 0.1543

P-value = 0.3086‬

\alpha = 0.10

P-value > \alpha

Fail to reject H0.There is no sufficient evidence to conclude the original claim that the mean of the population of student course evaluations is equal to 4.25 mean.